A tugboat goes 100 miles upstream in 10 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current.
The speed of the tugboat is__mph, and the speed of the current is__mph.

Answer by mananth(16946)

Tugboat speed =x mph
current speed =y mph
against current 10 hours
with current 5 hours

Distance against 100 miles distance with wind 100 miles
t=d/r against wind -
100 / ( x y y )= 10
10 ( x - y ) = 100
10 x - 10 y = 100 ....................1

100 / ( x y y )= 5
5 ( x + y ) = 100
5 x + 5 y = 100 ...............2
Multiply (1) by 1
Multiply (2) by 2
we get y
10 x + -10 y = 100
10 x + 10 = 200
20 x = 300
/ 20
x = 15 mph

plug value of x in (1) y
10 x -10 y = 100
150 -10 -150 = 100
-10 y = 100
-10 y = -50 mph
y = 5
Tugboat speed 15 mph
current speed 5 mph

m.ananth@hotmail.ca